### 25 Apr 2026 v3

# LinguDistill : Recovering Linguistic Ability in Vision-Language Models via Selective Cross-Modal Distillation

Abstract
Adapting pretrained language models (LMs) into vision-language models (VLMs) can degrade their native linguistic capability due to representation shift and cross-modal interference introduced during multimodal adaptation. Such loss is difficult to recover, even with targeted task-specific fine-tuning using standard objectives. Prior recovery approaches typically introduce additional modules that act as intermediate alignment layers to maintain or isolate modality-specific subspaces, which increases architectural complexity, adds parameters at inference time, and limits flexibility across models and settings. We propose LinguDistill , an adapter-free distillation method that restores linguistic capability by utilizing the original frozen LM as a teacher. We overcome the key challenge of enabling vision-conditioned teacher supervision by introducing layer-wise KV-cache sharing, which exposes the teacher to the student’s multimodal representations without modifying the architecture of either model. We then selectively distill the teacher’s strong linguistic signal on language-intensive data to recover language capability, while preserving the student’s visual grounding on multimodal tasks. As a result, LinguDistill recovers $\sim$10% of the performance lost on language and knowledge benchmarks, while maintaining comparable performance on vision-heavy tasks. Subject to further scaling tests, our findings suggest that linguistic capability of multimodal models can be recovered without extra modules, providing an efficient solution to modality degradation.
1 Introduction
Building multimodal models requires aligning different modalities into a shared representation space. In vision-language models (VLMs), this is typically achieved through fusion-based architectures (Liu et al., 2023a; Bai et al., 2025; Li and Tang, 2024; Zhao et al., 2025), which train the backbone language model (LM) to understand non-text modalities, or through earlier bridging paradigms that keep the LM frozen (Li et al., 2023; Alayrac et al., 2022). While the former generally demonstrates stronger performance, it comes at a cost. Adapting pretrained LM to multimodal settings leads to degradation in their native language capabilities, even pronounced when evaluated on purely textual inputs (Srivastava et al., 2024; Wang et al., 2025). This phenomenon, often attributed to modality interference or catastrophic forgetting, suggests that multimodal adaptation perturbs the linguistic priors encoded in the original backbone.
A natural approach is to further fine-tune the model on language-heavy tasks. However, prior work suggests that this is insufficient to fully recover language capability, as multimodal adaptation causes a representation shift and cross-modal interference that are not easily reversed through lightweight fine-tuning (Wang et al., 2025). As a result, language ability remains degraded even after further adaptation. To address this, prior work has explored several alternative directions. Bridging approaches, such as BLIP-style alignment, keep the LM largely frozen and introduce intermediate alignment modules to preserve its linguistic priors (Wang et al., 2024a; b; Zhang et al., 2023). Alternatively, subspace or weight-space methods attempt to recover lost capabilities by manipulating model parameters, for example through weight interpolation, task arithmetic, or selective parameter restoration (Wortsman et al., 2022; Ilharco et al., 2023; Zhu et al., 2024; Zhang and others, 2024). However, these approaches either rely on additional alignment modules, require large-scale training from scratch, or assume that desired capabilities can be cleanly isolated and restored in parameter space, which may not hold across models and modalities.
In this paper, we propose LinguDistill , a cross-modal knowledge distillation method to restore language capability in VLMs. We treat the frozen pretrained LM as a teacher and the VLM as a student, which is optimized using a standard distillation objective. A key challenge is that the teacher only operates on text, so its outputs are not aligned with vision-conditioned generation. To address this, we introduce a KV-cache sharing mechanism between the teacher and student decoders. This allows the LM to access the same multimodal context as the VLM and produce vision-aware supervision signals. After training, the teacher is removed, resulting in a standard VLM with improved language capability and no additional parameters or inference cost.
Our contributions are as below:
  • We propose an adapter-free selective distillation method that recovers native language capability in VLM by using the frozen original LM backbone as a teacher.
  • We propose a KV sharing training architecture to enable frozen LM to attend to multimodal context with no extra modules.
  • Our findings show that LinguDistill recovers performance on language/knowledge-oriented benchmarks by $\sim$10%, while exhibiting comparable degradation in vision-heavy tasks relative to standard fine-tuning methods.
2 Related Work
Preserving Language Capability in Multimodal Models.
Fine-tuning a pretrained language model on multimodal data degrades its language capability, a form of catastrophic forgetting documented in several recent studies (Zhai et al., 2023; Srivastava et al., 2024). The dominant approach to this problem freezes the language backbone entirely and inserts learnable modules to handle visual input. Frozen (Tsimpoukelli et al., 2021) trains only a vision encoder whose outputs are fed to a fully frozen GPT-J. BLIP-2 (Li et al., 2023) adds a Q-Former between image features and the frozen LM. CogVLM (Wang et al., 2024b) injects visual experts as parallel attention branches at every transformer layer. LLaMA-Adapter (Zhang et al., 2023) adds zero-initialised attention adapters to the top layers. IAA (Wang et al., 2024a) inserts inner adapters that alternate visual and textual processing at varying depths. These methods preserve language capability by never updating the LM, but they all add parameters that increase inference cost. Our method instead recovers language capability after standard fine-tuning, using the original LM as a distillation teacher and adding no parameters to the final model.
Cross-Modal Knowledge Distillation.
Knowledge distillation (Hinton et al., 2015) trains a student to match the soft output distribution of a teacher via KL divergence, combined with a hard cross-entropy loss on ground-truth labels. Cross-modal KD extends this to settings where teacher and student operate on different input modalities. Xue et al. (2023) show that cross-modal KD works because the student learns modality-general features shared across modalities. C2KD (Huo et al., 2024) provides a general framework for bridging the modality gap between teacher and student across audio-visual, image-text, and RGB-depth pairs. In the LLM era, Wang et al. (2025) distil a text-only LLM into a speech LLM, transferring linguistic capability across modalities. In vision-language settings, existing KD work focuses on model compression, transferring knowledge from a large VLM to a smaller one. In contrast, we distil from a text-only LM into a multimodal VLM, to recover linguistic capability lost during fine-tuning.
KV-Cache Sharing across Model Components.
Sharing key-value representations between model components lets one model attend to another’s context without re-encoding. Speculative decoding (Leviathan et al., 2023) lets a verifier extend the draft model’s KV cache. In encoder-decoder architectures, cross-attention over the encoder’s KV cache is the standard mechanism for the decoder to access input representations (Vaswani et al., 2017). T5Gemma 2 (Zhang et al., 2025) concatenates encoder and decoder KV representations into a single attention module with shared parameters. UniFusion (Li et al., 2025) explores layerwise KV fusion between a VLM encoder and a diffusion decoder, though they find naive concatenation without dedicated projections causes feature misalignment. We extend the student VLM’s KV cache to a frozen LM during training, allowing it to attend to visual context without modifying its parameters, and discard it after training.
3 LinguDistill
LinguDistill is a training-time framework that combines architectural modification and selective distillation to recover the native linguistic capability of a VLM. The required components are a VLM and its instruction-tuned LM backbone prior to multimodal alignment, such as LLaVA and its Vicuna backbone (Liu et al., 2023a). The VLM serves as the student and the frozen LM acts as the teacher whose linguistic capability is meant to be distilled back to the student. After training, the teacher is removed, resulting in a standard VLM with no additional parameters or inference overhead, which we denote as LinguDistill for the remainder of the paper.
3.1 KV Sharing Architecture
To enable frozen LM to have multimodal understanding, we leverage the VLM as the direct bridge. By taking inspiration from UniFusion’s layer-wise attention fusion (Li et al., 2025), we design layer-wise KV cache sharing. This is straightforward since the teacher and student share the same LM architecture. This allows the teacher to access the same multimodal representation produced by student for decoding context. During training, the student learns to produce multimodal representations that can be interpreted by the frozen LM so it can generate vision-aware supervision signals.
We first process the image using a vision encoder followed by a projector, mapping visual features into the language embedding space. These features are then combined with text tokens to form a multimodal sequence $\mathbf{X}=[\mathbf{X}_{v};\mathbf{X}_{t}]$. This multimodal representation can be obtained either through a standard encoder–projector pipeline or via unified paradigms (e.g., discrete tokenization or codebook-based representations). Ultimately, all cross-modal interaction happens at the decoder level. We denote the student decoder as $\Phi$ and the frozen teacher decoder as $\Omega$. The KV-cache sharing mechanism proceeds as follows:
  1. 1.
    The student decoder $\Phi$ processes the full multimodal sequence $\mathbf{X}$ and produces KV caches:
    $\mathbf{K}_{\Phi}^{(l)},\mathbf{V}_{\Phi}^{(l)}=\mathbf{X}\mathbf{W}_{k,\Phi}^{(l)},\;\mathbf{X}\mathbf{W}_{v,\Phi}^{(l)},\quad l=1,\dots,L$
  2. 2.
    The teacher decoder $\Omega$ reprocesses the same text prompt to form its query states:
    $\mathbf{Q}_{\Omega}^{(l)}=\mathbf{X}\mathbf{W}_{q,\Omega}^{(l)},\quad l=1,\dots,L$
  3. 3.
    The teacher $\Omega$ directly reuses the student KV cache at every layer:
    $\mathbf{K}_{*}^{(l)}=\mathbf{K}_{\Phi}^{(l)},\quad\mathbf{V}_{*}^{(l)}=\mathbf{V}_{\Phi}^{(l)}$
  4. 4.
    The teacher attends using its computed queries over the transported student memory:
    $\mathbf{A}^{(l)}=\mathrm{Softmax}\!\left(\frac{\mathbf{Q}_{\Omega}^{(l)}(\mathbf{K}_{\Phi}^{(l)})^{T}}{\sqrt{d_{h}}}+\mathbf{M}\right)\mathbf{V}_{\Phi}^{(l)}$
This design allows the teacher to attend to the student’s multimodal context via shared KV caches, effectively conditioning it on multimodal input. The approach is general and extends beyond bimodal interaction as well.
3.2 Selective Distillation Objective.
The student ($\Phi$) is optimized with a typical KD mixture of frozen LM’s ($\Omega$) soft distillation and its own hard next-token supervision. Let $\Omega_{\text{pos}}$ denote the set of non-padding positions of teacher models’ output whose labels are not ignored. With temperature $T$, the soft distillation loss is
$\mathcal{L}_{\mathrm{soft}}=\frac{T^{2}}{|\Omega_{\text{pos}}|}\sum_{(b,t)\in\Omega_{\text{pos}}}\mathrm{KL}\!\left(\mathrm{softmax}(\mathbf{z}^{\Omega}_{b,t}/T)\,\|\,\mathrm{softmax}(\mathbf{z}^{\Phi}_{b,t}/T)\right)$
We also retain the standard hard-label objective on the student:
$\mathcal{L}_{\mathrm{hard}}=\frac{1}{|\Omega_{\text{pos}}|}\sum_{(b,t)\in\Omega_{\text{pos}}}\mathrm{CE}(\mathbf{z}^{\Phi}_{b,t},y_{b,t})$
Instead of applying uniform distillation, LinguDistill uses data-dependent weighting, where distillation is primarily applied to language-intensive data. The final objective is
$\mathcal{L}=\frac{1}{|\Omega_{\text{pos}}|}\sum_{(b,t)\in\Omega_{\text{pos}}}\alpha(d_{b})\,\mathcal{L}^{(b,t)}_{\mathrm{soft}}+\left(1-\alpha(d_{b})\right)\,\mathcal{L}^{(b,t)}_{\mathrm{hard}}$
where $d_{b}$ denotes the data source of example $b$.
This objective allows the teacher’s supervision signal to be sourced more from language-relevant data, while relying more on hard supervision for visual-heavy sources. As a result, LinguDistill restores linguistic capability by applying distillation only on language-capability recovering data, where the LM provides strong supervision, while preserving the original structure of the VLM on visual-heavy or general multimodal data.
4 Setup
4.1 Dataset Design
All experiments use The Cauldron (Laurençon et al., 2024), a collection of 50 VL instruction-tuning datasets. From there, we retain 17 core sources covering a balanced mix of VQA, OCR/document, knowledge, and domain-specific tasks. This selection reduces the total dataset size to under 600k examples while preserving the most relevant training signals. From these 17 sources, we further define a subset of 8 language-intensive VL datasets for data ablation. We refer to these two settings as full (17 sources) and lang-subset (8 sources) and the summary table can be observed in Table 1. The full set serves as the baseline, simulating continued multimodal fine-tuning and enabling selective distillation via source identifiers. In contrast, the lang-subset is used to test whether restricting training to language-heavy data alone can recover linguistic capability. All samples contain 1 image, with a maximum sequence length of 1024 tokens, with rationales detailed in Appendix D.
Table 1: Training source categorisation. Language-heavy sources receive the full KD signal ($\alpha>0$); OCR/document sources receive CE only ($\alpha=0$) in selective distillation variants.
Category Sources KD signal Reasoning
Language-heavy (8)
A-OKVQA, FigureQA, IconQA,
RobustQA, ScienceQA,
Visual7W, VQAv2, VSR
$\alpha>0$
Distill teacher’s linguistic
priorson reasoning
and knowledge tasks
OCR/doc-heavy (9)
Chart2Text, ChartQA, DocVQA,
InfographicVQA, OCR-VQA,
TextCaps, TextVQA,
VisText, VisualMRC
$\alpha=0$
Preserve student’s native
visual and OCR grounding
ability
4.2 Models & Optimization
We use the instruction-tuned nanoVLM-460M-8k model (Wiedmann et al., 2025) as our VLM, along with its instruction-tuned LM backbone, SmolLM-360M-Instruct (Allal et al., 2025), as the teacher. The VLM serves as the student, with vision encoder being frozen, while the LM remains frozen during training. Both models are capped at a maximum input length of 1024 tokens, aligned with the dataset design. For training, we use the selective distillation objective described in Sec. 3.2. All variants are trained with the same learning rate at 1e-4, bf16 precision, and token budget at $\sim$380M tokens to 4000 steps using cosine LR scheduler. Expanded training details are provided in the Appendix A and C.
4.3 Experiment Runs
We design 10 run variants which serve as a baseline, method verification, and further data and hyperparameter ablations. All variants share the same base model, lusxvr/nanoVLM-460M-8k (Wiedmann et al., 2025), and draw from the data pool explained in Sec. 4.2. Table 2 summarises the configuration of each variant.
Table 2: Experimental variants. Baselines (gray), uniform KD (yellow), and selective KD (blue).
Variant KD $\boldsymbol{\alpha}$ $\boldsymbol{T}$ Notes
Baselines
nanoVLM Pretrained checkpoint
nanoVLM-full $\times$ Fine-tune on all sources
nanoVLM-lang $\times$ Fine-tune on language subset
Uniform KD
distill-full $\checkmark$ 0.5 2 KD on full subset
distill-lang $\checkmark$ 0.5 2 KD on lang subset
Selective KD (Ours)
LinguDistill $\checkmark$ 0.7 4 Selective KD on full data
LinguDistill-highKD $\checkmark$ 0.7 4 Higher Teacher Involvement
LinguDistill-lowKD $\checkmark$ 0.3 2 Lower Teacher Involvement
The variants are organized into three groups. Baselines establish the effect of standard fine-tuning without distillation, comparing the pretrained nanoVLM-full checkpoint with full-data fine-tuning and a language-only subset (nanoVLM-full, nanoVLM-lang). These measure how linguistic capability degrades under multimodal adaptation and whether data restriction alone can mitigate it. Uniform KD introduces distillation from the frozen LM across all data (distill-full), testing whether the teacher can recover linguistic capability. While this improves language-heavy benchmarks, it consistently harms vision-heavy tasks due to misalignment between text-only supervision and visual grounding. Selective KD (Ours) resolves this trade-off by routing the distillation signal based on data type. LinguDistill-lang applies KD only on language-heavy data, while LinguDistill-full retains all sources but suppresses KD on OCR/document tasks, preserving visual alignment while recovering linguistic capability. Additional ablations on temperature and distillation weight confirm that the gains arise from selective routing rather than hyperparameter choice.
4.4 Evaluation
We evaluate all variants using lmms-eval (Zhang et al., 2024) in a zero-shot setting with batch size 1. Rather than reporting a single aggregate score, we organize benchmarks into three groups that reflect distinct capability axes, allowing us to triangulate where linguistic recovery helps, where it hurts, and where effects are mixed. Framework identifier specific detail can be observed in Appendix B.
Language and knowledge-heavy benchmarks comprise ARC Easy and Challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), ScienceQA (Lu et al., 2022), and COCO 2017 captioning (Chen et al., 2015). These tasks are primarily solved through linguistic reasoning and world knowledge, with visual content playing a secondary role. We expect LinguDistill to show the gains here, as the distilled linguistic priors of the original LM backbone are most directly beneficial for this suite.
Document, OCR, and vision-specific benchmarks comprise DocVQA (Mathew et al., 2021), InfographicVQA (Mathew et al., 2022), OCRBench (Liu et al., 2023b), and RealWorldQA (xAI, 2024). These tasks demand fine-grained visual perception and are less reliant on linguistic priors. We expect some degradation here relative to standard fine-tuning, as distillation pressure toward language capability may reduce understanding on vision-heavy signals.
General multimodal benchmarks comprise MMMU (Yue et al., 2024), MMStar (Chen et al., 2024), MME perception and cognition (Fu et al., 2023), and AI2D (Kembhavi et al., 2016). These benchmarks blend visual understanding with knowledge and reasoning, making them a useful middle ground for assessing whether LinguDistill preserves overall multimodal competence while recovering linguistic capability.
5 Results
Table 3 reports benchmark scores across all variants. We group benchmarks by task type and organise variants from left (baselines) to right (our best method). All variants are trained starting from lusxvr/nanoVLM-460M-8k pretrained vanilla checkpoint and are evaluated at step 4000.
Standard fine-tuning degrades language capability.
Comparing nanoVLM and nanoVLM-full after 4000 training steps, we observe consistent regression on language-intensive and text-only benchmarks. COCO captioning drops from 0.800 to 0.673 ($15.9\%$), MME cognition from 302 to 229 ($24.2\%$), and HellaSwag from 0.405 to 0.326 ($19.5\%$). ARC Easy and ARC Challenge also drop by $10.7\%$ and $13.4\%$ respectively. These are benchmarks where generating or understanding language is the primary challenge. Meanwhile, DocVQA and MMStar ($0.3\%$) are barely affected, confirming that the degradation is concentrated on linguistic capability rather than visual understanding, confirming that further representation shift can degrade native linguistic capability more.
Table 3: Benchmark results across variants. Values are absolute scores; colored percentages denote change vs. nanoVLM-full. Average rows are manually color-scaled for perceptual consistency: small changes are subtle, larger changes progressively emphasized. Selective KD (LinguDistill) achieves the strongest gains coupled with a controlled perception ability degradation.
Fine-tuning Uniform KD (Ours w/o Task Weighting) Selective KD (Ours)
Task nanoVLM nanoVLM-full nanoVLM-lang distill-full distill-lang LinguDistill
Language-heavy
AI2D 0.440 0.416 0.413 (-0.7%) 0.490 (+17.8%) 0.502 (+20.7%) 0.507 (+21.9%)
COCO 0.800 0.673 0.742 (+10.3%) 0.843 (+25.3%) 0.787 (+16.9%) 0.866 (+28.7%)
ScienceQA 0.590 0.592 0.581 (-1.9%) 0.650 (+9.8%) 0.679 (+14.7%) 0.676 (+14.2%)
ARC Easy 0.605 0.540 0.542 (+0.4%) 0.548 (+1.5%) 0.601 (+11.3%) 0.621 (+15.0%)
ARC Challenge 0.322 0.279 0.281 (+0.7%) 0.270 (-3.2%) 0.296 (+6.1%) 0.318 (+14.0%)
HellaSwag 0.405 0.326 0.329 (+0.9%) 0.345 (+5.8%) 0.376 (+15.3%) 0.394 (+20.9%)
Average 0.527 0.471 0.481(+2.1%) 0.524(+11.3%) 0.540(+14.6%) 0.564(+19.7%)
Document & OCR
DocVQA 0.769 0.767 0.766 (-0.1%) 0.640 (-16.6%) 0.709 (-7.6%) 0.740 (-3.5%)
InfographicVQA 0.357 0.282 0.272 (-3.5%) 0.290 (+2.8%) 0.318 (+12.8%) 0.330 (+17.0%)
OCRBench 0.760 0.726 0.722 (-0.6%) 0.452 (-37.7%) 0.510 (-29.7%) 0.600 (-17.4%)
Average 0.629 0.592 0.587(-0.8%) 0.461(-22.1%) 0.512(-13.5%) 0.557(-5.9%)
General multimodal
MME Cognition 302 229 261 (+14.0%) 241 (+5.2%) 308 (+34.5%) 300 (+31.0%)
MME Perception 1259 1027 1070 (+4.2%) 1046 (+1.8%) 1109 (+8.0%) 1173 (+14.2%)
MME Avg 780.5 628.0 665.5(+6.0%) 643.5(+2.5%) 708.5(+12.8%) 736.5(+17.3%)
RealWorldQA 0.523 0.500 0.488 (-2.4%) 0.450 (-10.0%) 0.494 (-1.2%) 0.490 (-2.0%)
MMMU 0.320 0.310 0.302 (-2.6%) 0.283 (-8.7%) 0.288 (-7.1%) 0.283 (-8.7%)
MMStar 0.360 0.359 0.358 (-0.3%) 0.350 (-2.5%) 0.355 (-1.1%) 0.349 (-2.8%)
Non-MME Avg 0.401 0.390 0.383(-1.8%) 0.361(-7.4%) 0.379(-2.8%) 0.374(-4.1%)
Training on language-heavy data alone does not help.
nanoVLM-lang restricts training to the 8 language-heavy sources, removing all OCR and document data. This partially improves COCO ($10.3\%$ relative to nanoVLM-full) but shows inconsistent gains on reasoning benchmarks. HellaSwag improves only marginally to 0.329 ($0.9\%$), ARC Easy to 0.542 ($0.4\%$), while ARC Challenge remains limited at 0.281 ($0.7\%$). MME cognition improves to 261 ($14.0\%$), indicating partial recovery. However, we also observe slight regressions on AI2D ($0.7\%$) and ScienceQA ($1.9\%$), which are themselves language- and knowledge-intensive benchmarks. These degradations are counterintuitive, as such tasks should benefit from restricting training to language-heavy data, and suggest that the underlying representation drift is not resolved by data filtering alone. This suggests that while filtering helps reduce some degradation, it is insufficient to fully correct the backbone drift introduced by multimodal fine-tuning.
Uniform distillation partially improves language performance but damages vision tasks.
distill-full applies the KD signal uniformly across all sources with $\alpha{=}0.5$ and T=2T{=}2. Language benchmarks improve substantially over the fine-tuning baseline: COCO rises to 0.843 ($25.3\%$), AI2D to 0.490 ($17.8\%$), and ScienceQA to 0.650 ($9.8\%$). Text-only reasoning benchmarks also improve modestly: HellaSwag reaches 0.345 ($5.8\%$), ARC Easy 0.548 ($1.5\%$), while ARC Challenge remains slightly degraded at 0.270 ($3.2\%$). However, OCR and document tasks are severely damaged: OCRBench drops from 0.726 to 0.452 ($37.7\%$) and DocVQA from 0.767 to 0.640 ($16.6\%$). This signals that supervision from a pure language teacher, even with KV-cache sharing, misguides the student on pixel-level tasks lacking visual grounding. We conclude that uniform distillation over-regularizes the model toward language behavior, suppressing modality-specific signals and leading to task-specific degradation.
Selective distillation achieves the best language gains with contained vision degradation.
LinguDistill trains on all 17 sources (full subset) but applies $\alpha{=}0$ to the 9 OCR/document sources, selectively absorbing KD signal only for language-heavy examples. This results in the strongest improvements across all language benchmarks: ScienceQA reaches 0.676 ($14.2\%$), AI2D 0.507 ($21.9\%$), COCO 0.866 ($28.7\%$), and MME cognition 300 ($31.0\%$). Text-only reasoning benchmarks also show consistent gains: HellaSwag improves to 0.394 ($20.9\%$), ARC Easy to 0.621 ($15.0\%$), and ARC Challenge to 0.318 ($14.0\%$). These benchmarks have no visual component and directly reflect linguistic capability.
Vision task degradation is substantially reduced compared to uniform distillation: DocVQA drops only $3.5\%$ (vs. $16.6\%$), while OCRBench drops $17.4\%$ (vs. $37.7\%$). InfographicVQA instead improves to 0.330 ($17.0\%$), indicating better balance between modalities. We further analyze OCRBench degradation and find that these tasks rely on visual grounding signals unavailable to the text-only teacher (Appendix E). MMMU remains degraded at $8.7\%$ across distillation variants; we attribute this to the single-image, 1024-token training constraint rather than the distillation objective itself (Appendix D).
Overall, selective distillation achieves a superior trade-off: it maximizes language gains while preserving visual capability significantly better than uniform distillation.
6 Discussion
6.1 What Does Distillation Recover?
We compare LinguDistill and nanoVLM-full one by one to see what the teacher actually transfers. On ScienceQA, LinguDistill fixes 854 examples that nanoVLM-full gets wrong, while losing only 498 that nanoVLM-full gets right (+8.39 percentage points net). All of nanoVLM-full’s mistakes are actual wrong answers, not formatting issues. The gains cover natural science (443), language science (215), and social science (196), spanning physics, biology, language arts, and commonsense reasoning. The frozen teacher still knows things that the fine-tuned student has forgotten, and LinguDistill brings the knowledge back.
On AI2D, LinguDistill wins 645 examples and loses 362 (+9.16 percentage points), mostly in food web reasoning and biology diagrams. These tasks need the model to understand how things relate to each other, not just read labels. InfoVQA follows the same pattern. Even though it is an OCR-heavy benchmark, LinguDistill improves by +4.70 percentage points (703 wins vs. 432 losses) because the questions ask about what the infographic means, not just what it says. The key insight is that distillation helps whenever a task needs knowledge, regardless of whether it also involves reading text. The frozen teacher provides exactly that.
6.2 Effect on Visual Text Tasks
We break OCRBench into its ten sub-tasks (Appendix E) to see where the drops come from. Basic text recognition barely changes (4-4/150 points). The model can still read printed text. Most of the drop is in Key Information Extraction (63-63) and Doc-oriented VQA (26-26), which together make up 71% of the total loss. Both tasks require the model to pick the right field from a structured document like a receipt or form, where many text fields sit close together. A stronger language prior makes the model worse at picking the right spot.
We look at 172 examples where nanoVLM-full gets the answer right but LinguDistill does not, and find three failure patterns. First, the model corrects nonsense strings into real words (e.g., PEAEC becomes peace), which fails when the text is not meant to be a real word. Second, the model reads from the right document but the wrong spot, returning a valid answer from a nearby field. Third, the model paraphrases instead of copying, rounding numbers and shortening names. All three come down to the same thing. The teacher’s language signal makes the model better at understanding text but worse at copying exactly what it sees. This only shows up in document extraction tasks and does not spread to other benchmarks.
6.3 Ablations
Data Subset
Data subset does not significantly affect the gains. As shown in Table 3, comparing the baseline and Uniform KD settings, there is no substantial difference between using the full and lang subsets. This suggests that even targeted cross-modal recovery cannot be achieved through data selection alone, but instead requires architectural or optimization-level intervention.
Selective Distillation
We attribute LinguDistill ’s success to selective distillation. As observed when comparing Uniform KD and LinguDistill in Table 3, selective distillation not only preserves OCR capability but also improves text-only benchmarks. We argue that this is possible because the teacher signal is applied only where it is reliable (language-heavy data), while being suppressed on vision-heavy tasks, preventing misleading supervision and preserving the student’s visual grounding.
Distillation Hyperparams
Figure 1 shows the optimization behavior behind this trade-off. In Figure 1(a), we compare the CE term of our three selective distillation variants: the main LinguDistill setting, a stronger KD configuration (High KD), and a weaker one (Low KD). The main LinguDistill setting reaches the lowest CE curve overall, while High KD stays slightly above it and Low KD remains the highest throughout training. This indicates that pushing the teacher signal too hard hurts the student’s direct supervision, but weakening the KD term too much is also suboptimal. In Figure 1(b), we compare the corresponding KD losses. High KD gains the lowest soft loss, Low KD remains the highest soft loss, and the main LinguDistill setting sits between them. The best benchmark trade-off comes from the middle setting, where we balance learning through soft loss.
(a) Hard loss across KD-strength variants.
(b) Soft loss across KD-strength variants.
Figure 1: Training-loss analysis for the three selective distillation variants. Left: the CE term is lowest for the main LinguDistill setting, with High KD slightly above it and Low KD clearly worse. Right: the Soft loss uses the combined logged objective, which mixes the teacher-weighted distillation signal with the VLM loss in a balanced manner.
7 Conclusion
We study the degradation of linguistic capability in VLMs caused by multimodal adaptation, and show that this loss is difficult to recover with standard fine-tuning. We propose LinguDistill , an adapter-free distillation framework that restores linguistic capability by leveraging a frozen LM teacher through layer-wise KV-cache sharing and selective distillation. Our results demonstrate that LinguDistill recovers performance on language and knowledge benchmarks while maintaining comparable performance on vision-heavy tasks, surpassing standard finetuning objective with zero additional parameters. These findings suggest that targeted cross-modal distillation provides a simple and effective approach to preserving backbone capabilities in multimodal systems without introducing additional modules.
Appendix A Training Configuration
Table 4: Shared training configuration across all variants.
Parameter Value
Vision encoder SigLIP2-B/16-512 (Tschannen et al., 2025) (frozen)
Language decoder SmolLM2-360M-Instruct (Allal et al., 2025)
Image tokens 64 ($4{\times}$ pixel shuffle)
Max images per example 1
Max sequence length 1024 tokens
Effective batch size 128
Precision Mixed Precision BF16
Optimizer AdamW
LR (projector) $1{\times}10^{-4}$
LR (language decoder) $1{\times}10^{-4}$
LR (vision encoder) 0 (frozen)
LR schedule Cosine decay
Warmup 3% of total steps
Gradient clipping Norm 1.0
Dataset The Cauldron (Laurençon et al., 2024) (filtered)
Quality filters Relevance, image corr., visual dep., formatting $\geq 1$
All experiments use nanoVLM-full (Wiedmann et al., 2025) with a total size of 460M parameters, which pairs a SigLIP2-B/16 vision encoder (Tschannen et al., 2025) with a SmolLM2-360M-Instruct language decoder (Allal et al., 2025). We initialise from the lusxvr/nanoVLM-460M-8k checkpoint and train on a filtered subset of The Cauldron (Laurençon et al., 2024). All runs use a single RTX A6000 PRO (96 GB).
Table 4 lists hyperparameters shared across all variants.
For distillation variants, the teacher is a frozen copy of the original SmolLM2-360M-Instruct backbone used in nanoVLM-full. The teacher receives zero learning rate throughout training. All variants are evaluated at step 4000 unless noted otherwise in the main text.
Appendix B Evaluation Subsets
We evaluate all models using the lmms-eval framework (Zhang et al., 2024) with a standardized zero-shot protocol. All evaluations are performed with a fixed batch size of 1 on a single GPU, and predictions are logged for further analysis.
The evaluation tasks are selected directly from lmms-eval and correspond to the benchmark groups described in Sec. 4.4. Specifically, we use:
  • Language and knowledge-heavy: arc_easy, arc_challenge, hellaswag, scienceqa, coco2017_cap_val
  • Document and OCR: ocrbench, docvqa_val, infovqa_val, realworldqa
  • General multimodal: ai2d, mmstar, mme, mmmu_val
Appendix C Optimization Pseudocode
Algorithm 1 LinguDistill optimization pseudocode
Goal: Optimize the left-side VLM parameters using a detached right-decoder generation built via continuation over a shared KV memory.
Require: Minibatch $\mathcal{B}=\{(x_{b},I_{b},m_{b},y_{b},d_{b})\}_{b=1}^{B}$, image encoder $E_{\mathrm{img}}$, projector $P$, left decoder $D_{\mathrm{left}}$, right decoder $D_{\mathrm{right}}$, left head $H_{\mathrm{left}}$, right head $H_{\mathrm{right}}$, temperature $T$, source weighting rule $\alpha(\cdot)$.
Ensure: Scalar loss $\mathcal{L}$ and updated left-side parameters $\theta_{\mathrm{left}}$.
1 Initialize $\mathcal{L}_{\mathrm{sum}}\leftarrow 0$ and $N\leftarrow 0$
2 Freeze $D_{\mathrm{right}}$ and $H_{\mathrm{right}}$
3 for $b\leftarrow 1$ to $B$ do
4    $\mathbf{v}_{b}\leftarrow E_{\mathrm{img}}(I_{b})$
5    $\mathbf{u}_{b}\leftarrow P(\mathbf{v}_{b})$
6    $\tilde{\mathbf{x}}_{b}\leftarrow\mathrm{concat}(x_{b},\mathbf{u}_{b})$
7    $(\mathbf{h}^{l}_{b},\mathrm{KV}^{l}_{b})\leftarrow D_{\mathrm{left}}(\tilde{\mathbf{x}}_{b},m_{b})$
8    $\mathbf{z}^{l}_{b}\leftarrow H_{\mathrm{left}}(\mathbf{h}^{l}_{b})$
9    $\mathbf{h}^{r}_{b}\leftarrow D_{\mathrm{right}}(\tilde{\mathbf{x}}_{b},m_{b};\mathrm{KV}^{l}_{b})$
10    $\mathbf{z}^{r}_{b}\leftarrow\mathrm{stopgrad}(H_{\mathrm{right}}(\mathbf{h}^{r}_{b}))$
11    $\alpha_{b}\leftarrow\alpha(d_{b})$
12    for each $t$ such that $m_{b,t}=1$ and $y_{b,t}\neq-100$ do
13     $p^{l}_{b,t}\leftarrow\mathrm{softmax}(\mathbf{z}^{l}_{b,t}/T)$
14     $p^{r}_{b,t}\leftarrow\mathrm{softmax}(\mathbf{z}^{r}_{b,t}/T)$
15     $\ell_{b,t}\leftarrow\alpha_{b}T^{2}\mathrm{KL}(p^{r}_{b,t}\,\|\,p^{l}_{b,t})+(1-\alpha_{b})\mathrm{CE}(\mathbf{z}^{l}_{b,t},y_{b,t})$
16     $\mathcal{L}_{\mathrm{sum}}\leftarrow\mathcal{L}_{\mathrm{sum}}+\ell_{b,t}$
17     $N\leftarrow N+1$
18    end for
19 end for
20 $\mathcal{L}\leftarrow\mathcal{L}_{\mathrm{sum}}/N$
21 Update $\theta_{\mathrm{left}}=\{E_{\mathrm{img}},P,D_{\mathrm{left}},H_{\mathrm{left}}\}$ using $\nabla_{\theta_{\mathrm{left}}}\mathcal{L}$
22 Return $\mathcal{L}$
Figure 2: Pseudo-code for LinguDistill . The image is first mapped into the language subspace, the left decoder absorbs the full multimodal prompt and materializes prompt KV memory, and the frozen right decoder continues from that inherited state to produce detached supervision. The optimization objective combines source-selective distillation with hard next-token supervision, while updating only the left-side parameters.
Appendix D Sequence Length Analysis
We train with a maximum sequence length of 1024 tokens and a single image per example. This section justifies these choices with token-length statistics from the training data and a discussion of compute constraints.
Training data token distributions.
Table 5 reports token-length statistics for each training source, computed on a sample of up to 1000 examples per source using the SmolLM2-360M-Instruct tokenizer. The vast majority of examples fit within 1024 tokens. Only robut_sqa (2.4%) and scienceqa (2.1%) exceed this threshold, and in both cases the truncation affects fewer than 3% of examples. OCR and document sources have mean lengths under 130 tokens, well within the budget. The overall mean across all sources is under 120 tokens, meaning the 1024-token limit provides ample headroom for most training examples.
Table 5: Token-length statistics per training source, sampled from The Cauldron and tokenized with SmolLM2-360M-Instruct. “Cat.” indicates language-heavy (lang) or OCR/document (ocr) sources. “%>{>}1024” is the fraction of examples exceeding 1024 tokens.
Source Cat. Mean Med. P95 Max %>{>}1024
aokvqa lang 62 60 100 116 0.0%
figureqa lang 212 187 307 321 0.0%
iconqa lang 37 35 56 664 0.0%
robut_sqa lang 180 102 425 1974 2.4%
scienceqa lang 493 118 549 15171 2.1%
visual7w lang 251 197 705 972 0.0%
vqav2 lang 79 58 192 673 0.0%
vsr lang 38 26 80 98 0.0%
chart2text ocr 111 88 225 329 0.0%
chartqa ocr 43 35 92 134 0.0%
docvqa ocr 107 90 265 357 0.0%
infographic_vqa ocr 110 97 234 310 0.0%
ocrvqa ocr 102 102 124 162 0.0%
textcaps ocr 22 21 30 35 0.0%
textvqa ocr 31 34 48 56 0.0%
vistext ocr 128 126 182 217 0.0%
visualmrc ocr 99 70 241 292 0.0%
Evaluation benchmarks.
Our evaluation suite spans both short-sequence benchmarks (MME perception and cognition, RealWorldQA, ARC, MMStar) and longer-sequence benchmarks (DocVQA, InfoVQA, OCRBench, MMMU). The 1024-token training limit does not directly constrain evaluation, since generation uses a separate maximum output length. However, benchmarks that require long-context document understanding (DocVQA, InfoVQA) may be indirectly affected if the model has not seen sufficiently long sequences during training. The residual regressions on these benchmarks under selective distillation ($3.8\%$ on DocVQA, $7.6\%$ on InfoVQA) are consistent with this explanation. MMMU additionally requires multi-image reasoning, which our single-image training constraint does not support.
Compute constraints.
All experiments run on a single RTX A6000 PRO (96 GB). With BF16 precision, an effective batch size of 128, and the dual-tower distillation architecture (student + frozen teacher), a sequence length of 1024 fits comfortably in memory. Given that fewer than 1% of training examples exceed 1024 tokens, we prioritize training throughput over marginal coverage. Future work at larger scale can explore extended sequence lengths to improve document-heavy benchmarks.
Appendix E Detailed OCR Analysis
This appendix breaks down the OCRBench results from Section 6.
E.1 Sub-task Breakdown
OCRBench (Liu et al., 2023b) has 1,000 samples across ten sub-tasks. We group them by how much the model needs to understand document layout to answer correctly (Table 6).
Table 6: OCRBench sub-task breakdown grouped by cognitive demand. LinguDistill vs. nanoVLM-full (460M, 4k steps). Text recognition (decoding-only) is largely preserved; the regression concentrates in extraction tasks requiring spatial grounding (bold).
Group Sub-task LinguDistill nanoVLM-full $\Delta$
Text Rec. Regular /50 47 48 -1
Irregular /50 40 44 -4
Artistic /50 46 45 +1
Subtotal /150 133 137 -4
Handwriting & Non-Semantic Handwriting /50 30 36 -6
Digit String /50 38 42 -4
Non-Semantic /50 39 46 -7
Handwritten Math /100 46 52 -6
Subtotal /250 153 176 -23
Scene Text VQA /200 140 150 -10
Key Information Extraction /200 84 147 -63
Doc-oriented VQA /200 90 116 -26
Total /1000 600 726 -126
Text Recognition (150 samples).
This group tests reading printed text in regular, irregular (warped), and artistic (stylised) fonts. The model just needs to read the characters, with no document layout involved. LinguDistill scores 133/150 vs. nanoVLM-full’s 137/150, a gap of only 4 points. Regular text drops by one point (47 vs. 48) and artistic text improves by one (46 vs. 45), showing that the language prior can help with slightly ambiguous text.
Handwriting & Non-Semantic (250 samples).
This group covers handwritten words, digit strings, random letter strings (e.g., PEAEC), and math expressions, all of which are visually ambiguous. A language prior can either help or hurt here. The group drops by 23-23 points (153/250 vs. 176/250), with the biggest single drop in non-semantic text (7-7). This is where the language prior hurts most, since the ground truth is not a real word by design.
Layout-dependent tasks (600 samples).
Scene Text VQA (200 samples) adds a reasoning step where the model reads text in natural scenes and answers questions about it. Key Information Extraction (200 samples) and Doc-oriented VQA (200 samples) need the model to find the right field in structured documents (receipts, forms, tables) where many text fields sit close together. Key Information Extraction and Doc-oriented VQA together make up 89 of the 126-point total drop (71%), even though they are only 40% of the benchmark. Overall, the three layout-dependent groups account for 99 of 126 points (79%), while the four reading-only groups contribute only 27. We see the same pattern in DocVQA, where the biggest losses are in table/list questions (109-109), layout questions (84-84), and free-text extraction (89-89).
E.2 Failure Modes
We look at the 172 examples where nanoVLM-full gets the answer right but LinguDistill does not, and find three patterns.
The model corrects text into real words.
The language prior turns what the model sees into plausible words, even when the original text is not a real word. This shows up most in Non-Semantic Text Recognition, where LinguDistill fails on 11 of 50 items (vs. 4 for nanoVLM-full). For example, PEAEC becomes peace, meLtiid becomes meltid (dropping a letter to form a near-word), and eeorGg collapses to gg. In handwriting recognition, the characters are harder to read, which gives the language prior more room to step in. beer becomes 6er (a digit replaces an ambiguous stroke), both becomes 6.4, and soul becomes saul. nanoVLM-full reads these correctly because it has a weaker language prior. LinguDistill fails on 20 of 50 handwriting items, but 11 of those are shared failures where nanoVLM-full also gets it wrong, so about half the handwriting drop (9 items) comes from distillation.
The model reads from the wrong spot.
In Key Information Extraction, the model often reads from a nearby but wrong part of the document. When the expected answer is a receipt date (02/02/2018), the model returns the store hours (mon–sun: 1000 hrs – 2200 hrs). When the expected date is 06/03/18, it returns a timestamp from the next field (08/03/18 18:04). When the expected total is 14.20, it returns a subtotal from a different row (13.40). The answers are real text from the document, just from the wrong place. The stronger language prior makes the model favor fluent answers over finding the right field.
The model paraphrases instead of copying.
The model rewrites what it reads instead of copying it exactly. Numbers lose precision (5.04 becomes 5, 25.9 becomes 25). Descriptions get shortened (21–49 years of age becomes 21–49 years). Values get swapped with nearby ones (\$506 million becomes 193,690 from a different table cell). Names get paraphrased (program staff becomes nutrition health director). These are not errors in finding the right spot. The model sees the right region but rephrases what it reads, which is consistent with the teacher being trained on abstractive text.
Appendix F Expanded & Ablation Results
F.1 Full Results
Table 7: Full benchmark results across all variants. Columns are grouped by training strategy.
Fine-tuning Uniform KD Selective KD
Task nanoVLM nanoVLM-full nanoVLM-lang distill-full distill-lang LinguDistill highKD lowKD
Language and knowledge-heavy
AI2D 0.440 0.416 0.413 0.490 0.502 0.507 0.492 0.511
COCO 2017 Captioning2017_cap 0.800 0.673 0.742 0.843 0.787 0.866 0.876 0.842
ScienceQA 0.590 0.592 0.581 0.650 0.679 0.676 0.676 0.670
ARC Easy 0.605 0.540 0.542 0.548 0.601 0.621 0.614 0.627
ARC Challenge 0.322 0.279 0.281 0.270 0.296 0.318 0.315 0.314
HellaSwag 0.405 0.326 0.329 0.345 0.376 0.394 0.392 0.393
Document and OCR
DocVQA_val 0.769 0.767 0.766 0.640 0.709 0.740 0.720 0.747
InfographicVQA_val 0.357 0.282 0.272 0.290 0.318 0.330 0.320 0.322
OCRBench 0.760 0.726 0.722 0.452 0.510 0.600 0.598 0.601
General multimodal
RealWorldQA 0.523 0.500 0.488 0.449 0.494 0.490 0.500 0.484
mme_cognition 302 229 261 241 308 300 247 265
mme_perception 1259 1027 1070 1046 1109 1173 1082 1163
MMMU_val 0.320 0.310 0.302 0.283 0.288 0.283 0.293 0.288
MMStar 0.360 0.359 0.358 0.350 0.355 0.349 0.348 0.356
F.2 Data Subset Ablation
Table 8: Effect of data subset selection across all tasks.
Task nanoVLM nanoVLM-full nanoVLM-lang
Language and knowledge-heavy
AI2D 0.44 0.416 0.413
COCO 2017 Captioning 0.80 0.673 0.742
ScienceQA 0.59 0.592 0.581
ARC Easy 0.605 0.540 0.542
HellaSwag 0.405 0.326 0.329
Document and OCR
DocVQA 0.769 0.767 0.766
InfographicVQA 0.357 0.282 0.272
OCRBench 0.760 0.726 0.722
General multimodal
RealWorldQA 0.523 0.500 0.488
MME Cognition 302 229 261
MME Perception 1259 1027 1070
MMMU 0.32 0.31 0.302
MMStar 0.36 0.359 0.358
F.3 Uniform vs Selective Distillation Ablation
Table 9: Uniform vs selective distillation across all tasks.
Task distill-full distill-lang LinguDistill
Language and knowledge-heavy
AI2D 0.490 0.502 0.507
COCO 2017 Captioning 0.843 0.787 0.866
ScienceQA 0.650 0.679 0.676
ARC Easy 0.548 0.601 0.621
HellaSwag 0.345 0.376 0.394
Document and OCR
DocVQA 0.640 0.709 0.740
InfographicVQA 0.290 0.318 0.330
OCRBench 0.452 0.510 0.600
General multimodal
RealWorldQA 0.450 0.494 0.490
MME Cognition 241 308 300
MME Perception 1046 1109 1173
MMMU 0.283 0.288 0.283
MMStar 0.350 0.355 0.349
F.4 Distillation Hyperparameters Ablation
Table 10: Effect of distillation strength across all tasks.
Task LinguDistill highKD lowKD
Language and knowledge-heavy
AI2D 0.507 0.492 0.511
COCO 2017 Captioning 0.866 0.876 0.842
ScienceQA 0.676 0.676 0.670
ARC Easy 0.621 0.614 0.627
HellaSwag 0.394 0.392 0.393
Document and OCR
DocVQA 0.740 0.720 0.747
InfographicVQA 0.330 0.320 0.322
OCRBench 0.600 0.598 0.601
General multimodal
RealWorldQA 0.490 0.500 0.484
MME Cognition 300 247 265
MME Perception 1173 1082 1163
MMMU 0.283 0.293 0.288
MMStar 0.349 0.348 0.356